OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 6y = 36?

OpenStudy (anonymous):

Solve for y. The slope is the coefficient on the x variable. The y-intercept is the non-variable term

OpenStudy (bahrom7893):

6y = -3x + 36 y = -x/2 + 6 Slope is -1/2

OpenStudy (bahrom7893):

y intercept is 6

OpenStudy (anonymous):

First, subtract 3x to get 6y = -3x + 36. Then, divide by 6 to get y = (-3/6)x +6. That simplifies to y = (-1/2)x + 6. Now it's in slope intercept form, which means that the y-intercept is 6 and the slope is -1/2.